Is (1,-2) a solution of 3x-4y=12?

Thank you

Answer by bucky(2189)

All you have to do is to recognize that (1,-2) is a way of saying that x = 1 and y = -2.
You are given the equation 3x - 4y = 12
(1,-2) is a solution to this equation if you substitute 1 for x and -2 for y in the equation and the result is 12 = 12.
In other words:
3x - 4y = 12
3(1) - 4(-2) = 12
3 + 8 = 12
11 = 12
Notice that the left side of the equation is not equal to the right side. Therefore, (1, -2) is not a solution to the equation.